OpenStudy (anonymous):

c(x)=1.6x+3600 and r(x)=1.8x. How many units must be produced to break solid even

OpenStudy (anonymous):

@welshfella

OpenStudy (amistre64):

what does it mean to break even?

OpenStudy (anonymous):

It usually means cost vs revenue scenario?

OpenStudy (amistre64):

it is a cost to revenue scenario yes. it is when cost is equal to revenue

OpenStudy (amistre64):

1.6x + 3600 = 1.8x

OpenStudy (anonymous):

Would I have to subtract

OpenStudy (amistre64):

you would have to use any valid algebra process ... subtraction would work

OpenStudy (amistre64):

we could also multiply both sides by 10 and divide them by 2 16x + 36000 = 18x 8x + 18000 = 9x and then subtract 8x